Equivalent Units of Production The Converting Department of Hopkinsville Company had 640 units in work in process at the beginning of the period, which were 60% complete. During the period, 13,600 units were completed and transferred to the Packing Department.

There were 720 units in process at the end of the period, which were 30% complete. Direct materials are placed into the process at the beginning of production. Determine the number of equivalent units of production with respect to direct materials and conversion costs. If an amount is zero, enter in "0".

Hopkinsville Company Number of Equivalent Units of Production Whole Units Direct Materials Equivalent Units Conversion Equivalent Units Inventory in process, beginning Started and completed Transferred to Packing Department Inventory in process, ending Total
Answer #1

Answer

Number of Equivalent Units of Production

Whole Units

Direct Materials Equivalent Units

Conversion Equivalent Units

Inventory in process, beginning

                     640

                                  -  

                            256

Started and completed

                12,960

                          12,960

                       12,960

Transferred to Packing Department

                13,600

                          12,960

                       13,216

Inventory in process, ending

                     720

                               720

                            216

Total

                        14,320

                                     13,680

                                 13,432

--Workings

Number of Equivalent Units of Production

Whole Units

Direct Materials Equivalent Units

Conversion Equivalent Units

Inventory in process, beginning

640

=640*0%

=640*40%

Started and completed

=13600-640

=12960*100%

=12960*100%

Transferred to Packing Department

                13,600

                          12,960

                       13,216

Inventory in process, ending

720

=720*100%

=720*30%

Total

                        14,320

                                     13,680

                                 13,432